what is the efficiency of a diesel engine that receive 214.1 J of energy from combustion and lose 185.3 J by heat to exhaust during one cycle?


1
Expert's answer
2022-05-22T14:38:39-0400
"\eta=\frac{Q_H-Q_L}{Q_H}\\
=\frac{214.1-185.3}{214.1}=0.135=13.5\%"